# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v2

@@ IV The exciton transport dynamics by multichains
In this section, we discuss the effects of the inter-chain coupling, the dimerization parameter, the number and length of the chain on the exciton transport dynamics by numerically solving the master equation Eq. (14) under an external driving, which are displayed in two subsections below. According to the analysis in the Sec. III, the nearly-zero-energy modes are dressed by cavity for strong exciton-cavity coupling, namely the polaritonic branches forming. In order to let the polaritons act as a dominated role in exciton transport, we set that the Rabi splitting is far stronger than the decay rates of the cavity and the TLSs, i.e., $\Omega_{R,k}\gg(\kappa,\gamma)$.
@@ IV.1 The effects of the inter-chain coupling and the dimerization parameter
Here, we investigate primarily the corresponding relation between the energy level crossings (anticrossings) dressed by cavity and exciton transport dynamics to explain the exciton transport mechanism. Based on the relation and the properties of exciton transport dynamics, the optimal parameter ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$, as well as the optimal coupling configurations are expected to be presented.
@@
Fig. 6 shows the dependence of the steady-state exciton transport dynamics of the multi-chain system without cavity on $\zeta$ and $\delta$ for the six chosen configurations. From the Eq. (24), it can be known that the outgoing exciton current $I_{\rm o}$ is directly proportional to the total occupation number $\Sigma_{1}^{L}p_{k,N}$ of the outgoing sites, thus only $I_{\rm o}$ is shown in following figures. Hereinafter $\Sigma_{1}^{L}p_{k,N}$ is denoted by $p_{N}$. One can see that the effects of the inter-chain coupling and the dimerization parameter on exciton transport are significantly different, in which the steady-state occupation number $p_{11}$ of site ”11” (The two ”1” denote respectively the 1st row and the 1st column in the array of the TLSs shown in Fig. 1. This representation is similar below, unless otherwise noted.), $p_{\rm bulk}$ of bulk sites (The ”bulk” denotes the all sites from the 2nd to the $(N-1)$th column), the input and outgoing exciton current $I_{\rm i,o}$, and the exciton transport efficiency $\eta$ with $\zeta$ are symmetric about $\zeta=0$ in $[-1,1]$, but with $\delta$ are not.
@@
For the six configurations, the descending sort of efficiency is successively TC-SSH ST, TC ST, TC TT, TC-SSH TT, SSH ST, SSH TT in $[-1,1]$ [Fig. 6(a5)]. For the TC-SSH ST, near $\zeta=0.3$ the sites of bulk [Fig. 6(a2)], input, and outgoing possess commonly relatively high occupation numbers of exciton, which support exciton transport on multichains. Therefore, the exciton transport efficiency, input and outgoing currents are the largest [Fig. 6(a3,4,5)]. On the contrary, for $\zeta=0$ and $\zeta\rightarrow\pm 1$, although the site 11 possesses more excitations [Fig. 6(a1)], the exciton occupation number $p_{N}$ at outgoing sites is too few to transfer to the drain, leading the low efficiency and exciton currents. Specially, the excitons are distributed mainly at bulk sites for $\zeta\rightarrow\pm 1$ [Fig. 6(a2)], which is beneficial to energy storage and can be used to design quantum battery quach2022superabsorption . Except for the SSH ST, the other configurations also have the property for $\zeta\rightarrow\pm 1$, as well as TC TT and TC-SSH TT also do near $\zeta=\pm 0.4$. However, for TC ST near $\zeta=\pm 0.35$ and ±$\pm 0.97$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ are relatively high, thus the ability both of energy transport and storage are strong. Near $\zeta=\pm 0.6$, although the exciton transport efficiency is large, exciton currents are relatively small. Thus it is not beneficial to the exciton transport.
@@
It has four configurations SSH ST, SSH TT, TC-SSH ST, TC-SSH TT containing SSH chain. Because of the different spectra between odd and even SSH chain, in Fig. 6(b1-5) there are two situations $N=8,~{}9$ about the steady-state dynamics with $\delta$ in the absence of cavity. It is obvious that the exciton currents and transport efficiency are the largest near $\delta=0.3$ for the configurations SSH ST and TC-SSH ST of $N=8$ than that in the other range of $\delta$ and for the other configurations [Fig. 6(b3-5)], which are good for effective exciton transport. However, the SSH TT of $N=9$ is the most beneficial to energy storage because of the high occupation number $p_{\rm bulk}$ of bulk sites [Fig. 6(b2)] as well as the lowest exciton currents and transport efficiency. As shown in Fig. 6(b5), the exciton transport efficiency is high for $\delta<0$, but the exciton currents approach to zero, thus there is no real exciton transport. The reason is that the site 11 can not be coherently excited by drive, namely the excitons can not be injected in multichains [Fig. 6(b1,2)]. When $\delta>0.35$, the exciton currents and transport efficiency decrease sharply with $\delta$ or too small, because the excitons are localized at the left-edge site so that the occupation number of the right-edge sites is small for the TC-SSH ST and TC-SSH TT of $N=8,9$ or approaches to zero for the others.
@@
Under the TLS and the TLS+cavity drive, the effect of the inter-chain coupling $\zeta$ on the steady-state dynamics of the multi-chain system with cavity is presented for different configurations in Fig. 7(a1-6) and 7(b1-6) respectively. Particularly, to analyse what role the cavity acts as in exciton transport, we plot the dependence of the occupation number $p_{\rm c}$ of photons on $\zeta$ and mark its inflection points by the vertical fine lines in Fig. 7(a5), which correspond to the locations marked by the vertical fine lines in Fig. 2 and 3. Comparing Fig. 2, 3 with Fig. 7(a5), one can discover that each inflection point has corresponding crossings or anticrossings with the same locations. Next, we mainly analyse the effect of cavity on the exciton currents $I_{\rm i,o}$ and transport efficiency $\eta$. It can be seen that the presence of cavity breaks the symmetry of $I_{\rm i,o}$ with respect to $\zeta=0$ for configurations TC ST, SSH ST, SSH TT, and TC-SSH ST [Fig. 7(a3,4)]. Comparing to Fig. 6, it can be seen that the exciton transport is suppressed by cavity at the range of $\zeta\neq 0$. For example, near $\zeta=-0.35$ $p_{\rm c}$ is high [Fig. 7(a5)], suppressing the exciton currents in the main channel with the multichains for TC ST [Fig. 7(a3,4)]. On the other hand, there are three configurations SSH ST, SSH TT, and TC-SSH TT for which the cavity induces significant increase of $I_{\rm i,o}$ and $\eta$ near $\zeta=0$. In the situation, the exciton transport is assisted by polaritons, in which $p_{\rm c}$ is high, leading the increase of exciton occupation number on the last column sites $p_{N}$ [Fig. 7(a4)]. In fact, $\zeta=0$ means that only one chain undertakes the exciton transport, in which polariton-assisted exciton transport is demonstrated wei2022cavity . Especially, it can be seen obviously that the peaks of $p_{\rm c}$ and $I_{\rm i,o}$ are at different locations. The reason is that at the peak of $p_{\rm c}$ there is not enough excitations to be injected in multichains, causing relatively low $I_{\rm i,o}$. Because of few excitations injected [Fig. 7(a1)] for TC ST, TC TT and few $p_{N}$ [Fig. 7(a4)] for TC-SSH ST at the locations of the peaks of $p_{\rm c}$ cavity has no effect on exciton transport near $\zeta=0$. As shown in Fig. 7(b5), adding the cavity drive $p_{\rm c}$ increase largely, but $I_{\rm i,o}$ and $\eta$ do not. Near $\zeta=0$, because of the transfer of excitations the occupation number $p_{11}$ of site 11 increases [Fig. 7(b1)], causing the decrease of $p_{\rm c}$ for SSH ST, SSH TT, TC-SSH ST, and TC-SSH TT [Fig. 7(b5)], suppressing the exciton transport of the main channel with porlaritons [Fig. 7(b3,4,6)]. However, one can see by comparing Fig. 7(b3,4) with Fig. 7(a3,4) that the increase of $p_{\rm c}$ suppresses significantly the exciton currents of the main channel with the multichains near $\zeta=0.2$ for TC-SSH ST and near $\zeta=0.35$ for TC ST. In a word, the competition between cavity photons and excitons is the key affecting exciton transport.
@@
Fig. 8(a1-6) and 8(b1-6) show the dependence of the steady-state dynamics on the dimerization parameter $\delta$ under the TLS and the TLS+cavity drive in the presence of cavity respectively. We analyse firstly the effects of $\delta$ and cavity on exciton transport under the TLS drive. Although the exciton transport efficiencies are centered approximatively on $\delta=0$ [Fig. 8(a6)], there are no the exciton currents in $\delta<0$ except for the configurations SSH ST and TC-SSH ST of $N=9$ [Fig. 8(a3,4)]. The root cause is that excitons can not enter to multichains through the site 11 [Fig. 8(a1,3)]. It can be seen that for the configurations SSH ST and TC-SSH ST of $N=8$ in the range $0.350<\delta<$ 0.35 the presence of cavity suppresses significantly the exciton currents $I_{\rm i,o}$ [Fig. 8(a3,4)] and transport efficiency $\eta$ [Fig. 8 (a6)] by comparing to Fig. 6(b3-5). Particularly, at $\delta=0.26$ for SSH ST and $\delta=0.36$ for TC-SSH ST of $N=8$, the troughs of $I_{\rm i,o}$ and $\eta$ appear, which correspond to the peak and the trough of $p_{\rm c}$ [Fig. 8(a5)] and to the crossings near zero-energy modes dressed [Fig. 4] marked by the vertical fine lines, respectively. The reason for the former is that the excitation on the first and the last column TLSs is transferred to cavity through coherent exciting at cavity dressed crossings as shown in Fig. 8(a1,3,4), suppressing the exciton currents and transport efficiency on the main channel with multichains. But for the latter, the occupation number $p_{N}$ of exciton on the last column TLSs is too few, limiting the exciton transport. At $\delta>0.26$ for SSH ST and $\delta>0.36$ for TC-SSH ST of $N=8$, the limited photon number $p_{\rm c}$ [Fig. 8(a5)] and the decreased occupation number $p_{\rm bulk}$ on bulk sites [Fig. 8 (a2)] lead to that the polariton becomes the main channel, thus the exciton transport is strengthened by cavity comparing with Fig. 6(b3,4). However, for TC-SSH TT both the multichains and the polariton dominate the exciton transport because of the high $p_{\rm bulk}$ and increasing $p_{\rm c}$ with $\delta$. For the odd chains, in about $\delta>0.35$ the low $p_{N}$ is main reason limiting exciton transport. When the cavity drive is added based on the TLS drive, there are distinct effects on exciton transport in $0.350<\delta<$ 0.35 for SSH ST and TC-SSH ST of $N=8$. The increased $p_{\rm c}$ induces the increase of polaritons [Fig. 8(b5)], enhancing the exciton transport [Fig. 8(b3,4)] than that under TLS drive [Fig. 8(a3,4)], but in $\delta>0.35$ because of the low $p_{\rm c}$ and $p_{\rm bulk}$ [Fig. 8(b2)] the exciton transport is limited for SSH ST. Interestingly, the exciton currents and transport efficiency are significantly large in $\delta<0$ for SSH ST and TC-SSH ST of $N=9$, which break the universal situations that the exciton currents and transport efficiency are higher in $\delta>0$ than that in $\delta<0$. Specially, when the cavity is driven, the occupation number of site 11 is not necessary to transport exciton efficiently [Fig. 8(b1)]. In the condition, the high $p_{\rm bulk}$ [Fig. 8(b2)] and $p_{\rm c}$ [Fig. 8(b5)] illustrate that both the multichains and the polaritons participate the exciton transport. Although $p_{11}$ no longer is the condition restricting exciton transport under the TLS+cavity drive, $p_{N}$ still is. Therefore, the low $p_{N}$ suppresses the exciton transport for SSH TT and TC-SSH TT of $N=9$ in $\delta<0$.
@@
When the TLS transition frequency detuning $\Delta$ is considered, we plot the steady-state dynamics of the hybrid system in Fig. 9 and 10, which show respectively the effects of inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ for different configurations. Comparing with Fig. 2 and 3, the structures of exciton transport efficiency are similar to those of spectra. In Fig. 9 the zones of $\zeta$ and $\Delta$ that cavity has obviously effect on exciton transport are marked in Fig. 9(ab-fb, g-l, and gc-lc) by ellipses through comparing to Fig. 9(a-f). According to the effect these configurations can be classified into two categories. One is TC ST and TC TT that the exciton transport is suppressed, the other is the remaining configurations that the exciton transport is enhanced. For the former the transfer of excitation to cavity leads to the exciton transport channel with the multichains suppressed [Fig. 9(gc,hc)]. For the letter, as shown in Fig. 9(cb-fb) the occupation number $p_{\rm bulk}$ of bulk state is too small in the zone to support exciton transport of the channel with the multichains, thus the polaritons dominate the transport of exciton, in which the occupation number $p_{\rm c}$ of cavity photons is relatively large [Fig. 9(ic-lc)]. Further, we can illustrate the mechanism through spectra. Interestingly, the structure of $p_{\rm c}$ [Fig. 9(gc-lc)] is similar to that of the cavity-dressed energy levels [Fig. 2 and 3]. In the intensified zone by cavity there are the cavity-dressed anticrossings [Fig. 2 and 3], thus LZ transition is the foundation that cavity effects exciton transport. Particularly, in the high efficient zone of exciton transport there are two anti-crossed zonate zones with large $p_{\rm c}$ [Fig. 9(ic)] corresponding to the dressed anticrossings for SSH ST. When the cavity drive is exerted further, for the two categories of configurations the exciton transport is further suppressed and enhanced by cavity respectively. For SSH ST the effect of cavity drive on exciton transport is the biggest, in which the high efficient zone is significantly extended. In parameter space of $\delta$ and $\Delta$, as shown in Fig. 10 the steady-state dynamics depend on the coupling configuration, the cavity, and the method of drive. For SSH ST, TC-SSH ST of $N=8,~{}9$ and SSH TT of $N=8$, the cavity has apparent effects on exciton transport efficiency in the zone of the fully dimerized limit, namely, $\delta\rightarrow\pm 1$, and near the zero detuning [Fig. 10(a-d, g-i, m-p)]. Nevertheless, because the TC chain and the diagonal inter-chain coupling suppress the dimerized effect, leading to effect of cavity suppressed for TC-SSH TT [Fig. 10(s-v)]. Undoubtedly, in the zone enhancing exciton transport the polaritons act as prime role whose mechanism is coherent excitation (LZ transition) located at the cavity-dressed crossings (anticrossings) according to the analysis on Fig. 4, 5 and 8. When cavity drive is added, it can be seen that the high efficient zones are extended and the exciton transport efficiencies are strengthened in the zone of the fully dimerized limit and near the zero detuning for SSH ST, SSH TT, and TC-SSH ST of $N=8,~{}9$ [Fig. 10(e, f, k, l, q, r)].

### v3

@@ IV The exciton transport dynamics by multichains
In this section, we discuss the effects of the inter-chain coupling, the dimerization parameter, the number and length of the chain on the exciton transport dynamics by numerically solving the master equation Eq. (13) under an external driving, which are displayed in two subsections below. According to the analysis in the Sec. III, the nearly-zero-energy modes are dressed by the cavity for strong exciton-cavity coupling, namely the polaritonic branches forming. In order to let the polaritons act as a dominated role in exciton transport, we set that the Rabi splitting is far stronger than the decay rates of the cavity and the TLSs, i.e., $\Omega_{R,k}\gg(\kappa,\gamma)$.
@@ IV.1 The effects of the inter-chain coupling and the dimerization parameter
Here, we investigate primarily the corresponding relation between the energy level crossings (anticrossings) dressed by the cavity and the exciton transport dynamics to explain the exciton transport mechanism. Based on the relation and the properties of exciton transport dynamics, the optimal parameter ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$, as well as the optimal coupling configurations are expected to be presented.
@@
Fig. 6 shows the dependence of the steady-state exciton transport dynamics of the multi-chain system without cavity on $\zeta$ and $\delta$ for the six chosen configurations. From the Eq. (23), it can be known that the outgoing exciton current $I_{\rm o}$ is directly proportional to the total occupation number $\Sigma_{1}^{L}p_{k,N}$ of the outgoing sites, thus only $I_{\rm o}$ is shown in following figures, and $\Sigma_{1}^{L}p_{k,N}$ is denoted by $p_{N}$. One can see that the effects of the inter-chain coupling and the dimerization parameter on exciton transport are significantly different. The steady-state occupation number $p_{11}$ of site ”11” (The two ”1” denote respectively the 1st row and the 1st column in the array of TLSs shown in Fig. 1. This representation is similar below, unless otherwise noted.), $p_{\rm bulk}$ of bulk sites (The ”bulk” contains the all sites from the 2nd to the $(N-1)$th column), the input and outgoing exciton current $I_{\rm i,o}$, and the exciton transport efficiency $\eta$ with $\zeta$ are symmetric about $\zeta=0$ in $[-1,1]$, but with $\delta$ are not.
@@
For the six configurations, the descending sort of efficiency is successively HO-HE ST, HO ST, HO TT, HO-HE TT, HE ST, HE TT in $\zeta\in[-1,1]$ [Fig. 6(a5)]. For the HO-HE ST, near $\zeta=0.3$ the sites of bulk [Fig. 6(a2)], input, and outgoing possess commonly relatively high occupation numbers of exciton, which support exciton transport in multichains. Therefore, the exciton transport efficiency, input and outgoing currents are the largest near $\zeta=0.3$ for the HO-HE ST [Fig. 6(a3,4,5)]. On the contrary, for the HO-HE ST at $\zeta=0$ and $\zeta\rightarrow\pm 1$, although the site 11 possesses more excitations [Fig. 6(a1)], the exciton occupation number $p_{N}$ at outgoing sites is too few to transfer to the drain, leading the low efficiency and exciton currents. Specially, the excitons are distributed mainly at bulk sites for $\zeta\rightarrow\pm 1$ [see cyan line in Fig. 6(a2)], which is beneficial to energy storage and can be used to design quantum battery quach2022superabsorption . Except for the HE ST, the other configurations also have the property for $\zeta\rightarrow\pm 1$, as well as HO TT and HO-HE TT also do near $\zeta=\pm 0.4$. However, for HO ST near $\zeta=\pm 0.35$ and ±$\pm 0.97$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ are relatively high, thus the ability both of energy transport and storage are strong. Near $\zeta=\pm 0.7$ for HO ST, although the exciton transport efficiency is large, exciton currents are relatively small. Thus it is not beneficial to the exciton transport.
@@
It has four configurations HE ST, HE TT, HO-HE ST, HO-HE TT containing HE chain. Because of the different spectra between odd and even HE chain, in Fig. 6(b1-5) there are two situations $N=8,~{}9$ about the steady-state dynamics with $\delta$ in the absence of cavity. It is obvious that the exciton currents and transport efficiency are the largest near $\delta=0.3$ for the configurations HE ST and HO-HE ST of $N=8$ than that in the other range of $\delta$ and for the other configurations [Fig. 6(b3-5)], which are good for effective exciton transport. However, the HE TT of $N=9$ is the most beneficial to energy storage because of the high occupation number $p_{\rm bulk}$ of bulk sites [Fig. 6(b2)] as well as the lowest exciton currents and transport efficiency. Particularly, the exciton transport efficiency is high for $\delta<0$, but the exciton currents approach to zero, as shown in Fig. 6(b3-5), thus there is no real exciton transport. The reason is that the site 11 can not be coherently excited by drive, namely the excitons can not be injected in multichains [Fig. 6(b1,2)]. When $\delta>0.35$, the exciton currents and transport efficiency decrease sharply with $\delta$ or too small, because the excitons are localized at the left-edge site so that the occupation number of the right-edge sites is small for the HO-HE ST and HO-HE TT of $N=8,9$ or approaches to zero for the others.
@@
Under the TLS and the TLS+cavity drive, the effect of the inter-chain coupling $\zeta$ on the steady-state dynamics of the multi-chain system in a cavity is presented for different configurations in Fig. 7(a1-6) and 7(b1-6) respectively. Particularly, to analyse what role the cavity acts as in exciton transport, the dependence of the occupation number $p_{\rm c}$ of photons on $\zeta$ is shown, and its extreme points are marked by the vertical fine lines in Fig. 7(a5), which correspond to the locations of the crossings and anticrossings in spectra, as well as of the incontinuity and extreme points of the second order partial derivative $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of the photonic Hopfield coefficients, marked by the vertical fine lines in Fig. 2 and 3. In consequence, the second order partial derivative of the photonic Hopfield coefficients can also describe the change of the occupation number of photons. It can be seen that the presence of the cavity breaks the symmetry of $I_{\rm i,o}$ with respect to $\zeta=0$ for configurations HO ST, HE ST, HE TT, and HO-HE ST [Fig. 7(a3,4)]. Comparing with Fig. 6, one can discover that the exciton transport is suppressed by the cavity in the range of $\zeta\neq 0$. For example, near $\zeta=-0.3$ $p_{\rm c}$ is high [Fig. 7(a5)], suppressing the exciton currents in the main channel with the multichains for HO ST [Fig. 7(a3,4)]. On the other hand, there are three configurations HE ST, HE TT, and HO-HE TT for which the cavity induces significant increase of $I_{\rm i,o}$ and $\eta$ near $\zeta=0$. In the situation, the exciton transport is assisted by polaritons, in which $p_{\rm c}$ is high, leading the increase of exciton occupation number $p_{N}$ on the last column sites [Fig. 7(a4)]. In fact, $\zeta=0$ means that only one chain undertakes the exciton transport, in which polariton-assisted exciton transport is demonstrated wei2022cavity . Because of few excitations injected [Fig. 7(a1)] for HO ST, HO TT and few $p_{N}$ [Fig. 7(a4)] for HO-HE ST near $\zeta=0$, the cavity has no effect on exciton transport. What’s more, it can be seen obviously that the peaks of $p_{\rm c}$ and of $I_{\rm i,o}$ are at different locations. The reason is that at the peak of $p_{\rm c}$ the multichains do not occupy enough excitations when it dominates the exciton transport, causing relatively low $I_{\rm i,o}$. As shown in Fig. 7(b5), adding the cavity drive $p_{\rm c}$ increase largely, but $I_{\rm i,o}$ and $\eta$ do not. Near $\zeta=0$, because of the transfer of excitations the occupation number $p_{11}$ of site 11 increases [Fig. 7(b1)], causing the decrease of $p_{\rm c}$ for HE ST, HE TT, HO-HE ST, and HO-HE TT [Fig. 7(b5)], suppressing the exciton transport of the main channel with the porlaritons [Fig. 7(b3,4,6)]. However, one can see by comparing Fig. 7(b3,4) and Fig. 7(a3,4) that the increase of $p_{\rm c}$ suppresses significantly the exciton currents of the main channel with the multichains near $\zeta=0.2$ for HO-HE ST and near $\zeta=0.35$ for HO ST. Therefore, the exciton currents $I_{\rm i,o}$ and transport efficiency $\eta$ can be adjusted by the cavity via photons. According to different configurations, we have the conclusion that the coupling type ST is more favorable to exciton transport than TT, and the optimal locations of exciton transport for HO and HE are different in parameter space of $\zeta$, which are consistent with the spectra. In a word, the competition between cavity photons and excitons is the key affecting exciton transport.
@@
Fig. 8(a1-6) and 8(b1-6) show the dependence of the steady-state dynamics on the dimerization parameter $\delta$ under the TLS and the TLS+cavity drive in the presence of cavity respectively. We analyse firstly the effects of $\delta$ and cavity on exciton transport under the TLS drive. Although the exciton transport efficiencies are centered approximatively on $\delta=0$ [Fig. 8(a6)], there are no the exciton currents in $\delta<0$ except for the configurations HE ST and HO-HE ST of $N=9$ [Fig. 8(a3,4)]. The root cause is that the excitation can not enter to multichains through the site 11 [Fig. 8(a1,3)]. It can be seen that for the configurations HE ST and HO-HE ST of $N=8$ in the range $0.350<\delta<$ 0.35 the presence of cavity suppresses significantly the exciton currents $I_{\rm i,o}$ [Fig. 8(a3,4)] and transport efficiency $\eta$ [Fig. 8 (a6)] by comparing with Fig. 6(b3-5). Particularly, at $\delta=0.26$ for HE ST and $\delta=0.36$ for HO-HE ST of $N=8$, the troughs of $I_{\rm i,o}$ and $\eta$ appear, which correspond to the peak and the trough of $p_{\rm c}$ [Fig. 8(a5)] and to the crossings of the dressed near zero-energy modes marked by the vertical fine lines in Fig. 4, respectively. The reason for the former is that the excitation on the first and the last column TLSs is transferred to cavity through coherent exciting at the dressed crossings as shown in Fig. 8(a1,3,4), suppressing the exciton currents and transport efficiency on the main channel with multichains. But for the latter, the occupation number $p_{N}$ of exciton on the last column TLSs is too few, limiting the exciton transport. At $\delta>0.26$ for HE ST and $\delta>0.36$ for HO-HE ST of $N=8$, the limited photon number $p_{\rm c}$ [Fig. 8(a5)] and the decreased occupation number $p_{\rm bulk}$ on bulk sites [Fig. 8 (a2)] lead to that the polariton becomes the main channel, thus the exciton transport is strengthened by the cavity comparing with Fig. 6(b3,4). However, for HO-HE TT both the multichains and the polariton dominate the exciton transport because of the increase of $p_{\rm bulk}$ and $p_{\rm c}$ with $\delta$. For the odd chains, in about $\delta>0.35$ the low $p_{N}$ is main reason limiting exciton transport. When the cavity drive is added based on the TLS drive, there are distinct effects on exciton transport in $0.350<\delta<$ 0.35 for HE ST and HO-HE ST of $N=8$. The increased $p_{\rm c}$ [Fig. 8(b5)] induces the increase of polaritons, enhancing the exciton transport [Fig. 8(b3,4)] than that under TLS drive [Fig. 8(a3,4)], but in $\delta>0.35$ because of the low $p_{\rm c}$ and $p_{\rm bulk}$ [Fig. 8(b2)] the exciton transport is limited for HE ST of even chain. On the contrary, because of the high $p_{\rm c}$ and $p_{\rm bulk}$ the exciton transport is heightened for HO-HE ST and HO-HE TT of odd chain. Nevertheless, the high $p_{\rm bulk}$ and $p_{\rm c}$ make that the exciton currents and transport efficiency still are high in $\delta>0.7$ for HO-HE ST compared with that under TLS drive. Interestingly, the exciton currents and transport efficiency are significantly large due to the cavity drive in $\delta<0$ for all configurations of $N=9$, which break the universal situations that the exciton currents and transport efficiency are higher in $\delta>0$ than that in $\delta<0$. Specially, when the cavity is driven, the occupation number $p_{11}$ of site 11 is no longer necessary to transport efficiently exciton [Fig. 8(b1)]. In the condition, the high $p_{\rm bulk}$ [Fig. 8(b2)] and $p_{\rm c}$ [Fig. 8(b5)] illustrate that both the multichains and the polaritons participate the exciton transport. Although $p_{11}$ no longer is the condition restricting exciton transport under the TLS+cavity drive, $p_{N}$ still is. Therefore, the low $p_{N}$ limits the exciton transport for HE TT and HO-HE TT of $N=9$ in $\delta<0$. To sum up, in $\delta>0$ the cavity have significant influence to exciton transport, which the exciton transport is suppressed in weak dimerization strength and is enhanced in strong dimerization strength for even chain under TLS drive. In the space of $\delta$, the coupling type ST on even chain is more beneficial to exciton transport. However, when the cavity drive is exerted, ST on odd chain is good for exciton transport.
@@
When the TLS transition frequency detuning $\Delta$ is considered, we plot the steady-state dynamics of the hybrid system in Fig. 9 and 10, which show respectively the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ for different configurations. Comparing with Fig. 2 - 5, the structures of exciton transport efficiency are similar to those of spectra. In Fig. 9 the zones of $\zeta$ and $\Delta$ that cavity has obviously effect on exciton transport are marked in Fig. 9(ab-fb, g-l, and gc-lc) by ellipses through comparing to Fig. 9(a-f). According to the effect these configurations can be classified into two categories. One is HO ST and HO TT that the exciton transport is suppressed, the other is the remaining configurations that the exciton transport is enhanced. For the former the transfer of excitation to the cavity leads to the exciton transport channel with the multichains suppressed [Fig. 9(gc,hc)]. For the letter, as shown in Fig. 9(cb-fb) $p_{\rm bulk}$ is too small in the zone existing exciton transport in the absence of the cavity, thus the polaritons dominate the transport of exciton, in which the occupation number $p_{\rm c}$ of the cavity photons is relatively large [Fig. 9(ic-lc)]. Further, we can illustrate the mechanism through the spectra. Interestingly, the structures of $p_{\rm c}$ [Fig. 9(gc-lc)] are similar to those of the cavity-dressed energy levels [Fig. 2 and 3]. In the intensified zone by the cavity there are the cavity-dressed anticrossings [Fig. 2 and 3], thus LZ transition is the foundation that the cavity effects exciton transport. Particularly, in the high efficient zone of exciton transport there are two anti-crossed zonate zones with large $p_{\rm c}$ [Fig. 9(ic)] corresponding to the dressed anticrossings for HE ST. When the cavity drive is exerted further, for the two categories of configurations the exciton transport is further suppressed and enhanced by the cavity respectively. For HE ST the effect of cavity drive on exciton transport is the biggest than the other configurations, in which the high efficient zone is significantly extended. In parameter space of $\delta$ and $\Delta$, as shown in Fig. 10 the steady-state dynamics depend on the coupling configuration, the cavity, and the method of drive. For HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ the cavity has apparent effects on exciton transport efficiency in the zone of the fully dimerized limit, namely, $\delta\rightarrow\pm 1$, and near the zero detuning [Fig. 10(a-d, g-i, m-p)]. Nevertheless, the effect of the cavity is suppressed for HO-HE TT [Fig. 10(s-v)] because the HO chain and the diagonal inter-chain coupling suppress the dimerized effect. Undoubtedly, in the zone enhancing exciton transport the polaritons act as prime role whose mechanism is coherent excitation (LZ transition) located at the cavity-dressed crossings (anticrossings) according to the analysis on Fig. 4, 5 and 8. When the cavity drive is added, it can be seen that the high efficient zones are extended and the exciton transport efficiency is strengthened in the zone of the fully dimerized limit and near the zero detuning for HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ [Fig. 10(e, f, k, l, q, r)]. If the concern is only on the optimal efficiency of exciton transport, HE ST is the preferred configuration in the parameter space of $\zeta$ and $\Delta$ under the TLS drive in the presence of the cavity, as well as HE TT is the preferred configuration in the parameter space of $\delta$ and $\Delta$ under the TLS+cavity drive. Significantly, according to the results the exciton transport can be controlled by the coupling configurations, the cavity, the drive methods, the inter-chain coupling, the dimerization parameter, and the TLS transition frequency detuning.